OpenStudy (anonymous):

Solve the equation using square roots. A. –3, 3 B. –9, 9 C. D. no real number solutions

OpenStudy (anonymous):

x^2-9=0 is the problem

OpenStudy (anonymous):

.x^2-9=o add 9 on both sides x^2=9 square root both sides gives sqrt(x^2)=sqrt(9) x=?